# Criticality Leveraged Adversarial Training (CLAT) for Boosted Performance via Parameter Efficiency

### v2

@@ 4.3.1 Ablating on pretraining epochs before CLAT
As discussed in Section 3.3, we apply CLAT after the model has been adversarially trained for some epochs. Here, we analyze how the number of pretraining epochs affects CLAT performance. Figure 2 shows the training curves for different allocations of PGD pretraining epochs and CLAT fine-tuning epochs within a 100-epoch training budget. The overfitting of PGD-AT is evident as adversarial accuracy plateaus and declines towards the end, as documented in previous research [24]. In contrast, CLAT continues to improve adversarial accuracy, effectively addressing this issue. Including CLAT at any stage of training results in higher clean accuracy and robustness at convergence. Additional results on pretrained clean models are provided in Appendix C.
Furthermore, an intriguing aspect of our experiments involves running CLAT from scratch (0 PGD-AT epochs). Although CLAT ultimately surpasses PGD-AT with sufficient epochs, using CLAT without any prior adversarial training results in significantly slower model convergence. We believe this suggests that “layer criticality” emerges during the adversarial training process, allowing critical layers to be identified as the model undergoes adversarial training. This phenomenon supports our theoretical insight that criticality can be linked to the curvature of the local minima to which each layer converges during adversarial training. Flattening the local minima of critical layers aids in model generalization. We plan to explore this phenomenon further in future work.
Figure 2: Comparative analysis of CLAT performance on WRN34-10: Clean and adversarial accuracy on CIFAR-10 across partially trained models. Note: PGD-AT requires 150 epochs to fully converge.
@@
Table 4: Comparison of PGD-AT, PGD-AT + CLAT, and PGD-AT + CLAT (Non-dynamic) on CIFAR-10 Adv. Accuracy.
Method DN121 RN50
pgd-at 58.15 56.35
clat 60.60 59.54
clat (nd) 57.01 54.22
4.3.2 Ablating on critical layer selection
The choice of critical layer selection is another important feature impacting the performance of CLAT. We begin by examining the effect of dynamic layer selection. Table 4 highlights that dynamic selection is crucial to CLAT’s performance. Using the same layers throughout the process tends to cause overfitting and results in lower accuracies compared to the PGD-AT baseline.
To verify the significance of the selected critical layers, we compare CLAT with an alternative approach in which random layers are dynamically selected for fine-tuning instead of the critical layers. The results of this comparison are detailed in Table 6 and Table 10 (in Appendix C) on CIFAR-10 and CIFAR-100, respectively. The data demonstrates that selecting critical layers significantly enhances the model’s adversarial robustness and yields a marginal increase in clean accuracy. Moreover, Table 4 indicates near-identical critical layer selections within the same model even across diverse datasets. This evidence supports our assertion that the variation in layer criticality stems from inherent properties within the model architecture, wherein certain layers are predisposed to learning non-robust features.
Table 5: Critical vs. random layers for CLAT on CIFAR-10.
Network CIFAR-10
Critical Layers Random Layers
Clean Acc. Adv. Acc. Clean Acc. Adv. Acc.
DN121 81.03 60.60 78.85 51.35
RN50 83.78 59.54 79.01 51.44
RN18 83.89 55.37 78.02 51.03
@@ 4.3.2 Ablating on critical layer selection
vgg19 39.3M 236K 6.0%
Lastly, we ablate on the number of critical layers used in CLAT for fine-tuning. Figure 3 illustrates the trade-off between adversarial test accuracy and the quantity of critical layers selected for fine-tuning. Allowing more layers to be fine-tuned enhances the model flexibility, which initially improves CLAT performance. However, fine-tuning an excessive proportion of critical layers diminishes CLAT’s effectiveness. This is likely due to the diversion of attention towards fine-tuning less-critical layers, detracting from more critical ones. This pattern highlights the crucial influence of specific layers on network robustness and suggests a need for deeper research to understand the roles and dynamics of individual layers influencing network robustness.
Figure 3: Comparative analysis on CLAT performance/adversarial accuracy with respect to number of critical layers used during CLAT
4.4 Efficiency and stability analysis
To verify the significance of the selected critical layers, we contrast CLAT with an alternative approach in which random layers are dynamically selected for fine-tuning instead of the critical layers. The results of this comparison are detailed in Table 6 and Table 10 (in Appendix C) on CIFAR-10 and CIFAR-100, respectively.
The cost of optimizing the CLAT training objective in Equation 9 is similar to that of the standard adversarial training given its min-max formulation. Here, we show that the computational overhead for determining critical indices is negligible. We verify in Table 7 that criticality indices can be stably computed with a single training batch as small as 10, with top-ranking layers consistent with those achieved with a larger batch. We further conducted over 1000 runs for each network to randomly select the data used for criticality estimation, where we find remarkable consistency in computed critical layers, differing in less than 5% of cases, typically involving only one layer change among the top five critical layers. The stability means we can use a mere 0.0002% of the training data for criticality estimation every 10 epochs, which introduces neglectable additional complexity to the training process as indicated by the time measurement in the table.
@@ 5 Conclusions
In this work, we introduce CLAT, an innovative adversarial training approach that addresses adversarial overfitting issues by fine-tuning only the critical layers vulnerable to adversarial perturbations. This method not only emphasizes layer-specific interventions for enhanced network robustness but also sheds light on the commonality in non-robust features captured by these layers, offering a targeted and effective defense strategy. Our results reveal that CLAT reduces trainable parameters by about 95% while ensuring significant improvements in clean accuracy and adversarial robustness across diverse network architectures and baseline adversarial training methods. We limit the scope of this work to improving the empirical robustness of the model by utilizing the critical layers. In this sense, open questions remain on why these specific layers become critical, whether they can be identified more effectively, and whether the issues can be resolved with architectural or training scheme changes. We leave a more theoretical understanding of these questions as future work.
@@ Appendix A Pseudocode of CLAT
7:         Compute $\mathcal{W}_{\epsilon}(F_{i})$ for all layers with Equation 5
8:         Compute $\mathcal{C}_{f_{i}}$ for all layers with Equation 6
@@
13:         Perturbation $\delta\leftarrow$ Equation 8 inner maximization
14:         $\mathcal{L}_{C}(f_{\mathcal{S}})\leftarrow$ Equation 8
15:         Weight update $w[\mathcal{S}]$ with Equation 9
@@ Appendix B Additional experiment results
Table LABEL:wb_aa showcases results under a variant of the white-box Auto Attack [8] with an epsilon value of 0.03 and a single restart in Table 1. Specifically, we report adversarial accuracies of the “APGDCE” untargeted attack, a component of the Auto Attack ensemble, with similar trends observed for other attacks in the ensemble.
Although CLAT models employ PGD-like attacks on hidden features during the adversarial training phase and have not seen Auto Attacks, the robustness remains under Auto Attacks. This highlights the similarity in non-robust features utilized by different attacks in different networks. As CLAT resolves the non-robust features with critical layer fine-tuning, the resulting robustness is generalizable across attack settings and source models of the attacks.
Table 8: Adversarial accuracy under the white-box “APGDCE” variant of Auto Attack: Models trained using CLAT consistently outperform those trained exclusively with PGD-AT.
Dataset Network PGD-AT(%) PGD-AT + CLAT (%)
CIFAR-10 DN121 49.26 50.58
WRN70-16 54.48 56.90
RN50 48.54 50.65
WRN34-10 52.21 54.73
VGG19 45.35 47.01
RN18 46.99 48.20
CIFAR-100 DN121 24.67 26.31
WRN70-16 30.27 33.32
RN50 25.44 26.64
WRN34-10 27.62 31.53
VGG19 22.50 24.02
RN18 23.81 24.43
We further compare the CLAT model robustness with the robustness of the SAT model against white-box attacks of various strengths. As illustrated in  Figure 4, though both models are only trained against an attack of one strength ($\epsilon$ = 0.03), the improved robustness of CLAT is consistent across the full spectrum of attack strengths. This shows that CLAT is not overfitting to the specific attack strength used in training.
@@ C.2 Additional results on layer selection
Table 10: Ablating the layer choices for CLAT fine-tuning on CIFAR-100.
Network CIFAR-100
Critical Layers Random Layers
Clean Acc. Adv. Acc. Clean Acc. Adv. Acc.
DN121 58.79 33.23 54.85 23.32
RN50 61.88 36.23 56.50 22.45
RN18 59.22 32.04 52.43 22.60

### v3

@@ 4.3.1 Ablating on pretraining epochs before clat
As discussed in Section 3.3, we apply CLAT after the model has been adversarially trained for some epochs. Here, we analyze how the number of pretraining epochs affects CLAT performance. Fig. 2 shows the training curves for different allocations of PGD pretraining epochs and CLAT fine-tuning epochs within a 100-epoch training budget. The overfitting of PGD-AT is evident as adversarial accuracy plateaus and declines towards the end, as documented in previous research (Rice et al., 2020). In contrast, CLAT continues to improve adversarial accuracy, effectively addressing this issue. Including CLAT at any stage of training results in higher clean accuracy and robustness at convergence. Additional results on pretrained clean models are provided in Appendix C.
Furthermore, an intriguing aspect of our experiments involves running CLAT from scratch (0 PGD-AT epochs). Although CLAT ultimately surpasses PGD-AT with sufficient epochs, using CLAT without any prior adversarial training results in significantly slower model convergence. We believe this suggests that “layer criticality” emerges during the adversarial training process, allowing critical layers to be identified as the model undergoes adversarial training. This phenomenon supports our theoretical insight that criticality can be linked to the curvature of the local minima to which each layer converges during adversarial training.
Figure 2: Comparative analysis of CLAT performance on WRN34-10: Clean and adversarial accuracy on CIFAR-10 across partially trained models.
Table 5: Adversarial accuracy under PGD attack: Comparison of PGD-AT, PGD-AT + CLAT, and PGD-AT + CLAT (Non-dynamic) on CIFAR-10 and CIFAR-100.
Method CIFAR-10 CIFAR-100
DN121 RN50 DN121 RN50
pgd-at 58.15 56.35 31.76 33.01
clat 60.60 59.54 33.23 36.23
clat (nd) 57.01 54.22 30.34 32.98
Table 6: Adversarial accuracy under Auto attack: Comparison of PGD-AT, PGD-AT + CLAT, and PGD-AT + CLAT (Non-dynamic) on CIFAR-10 and CIFAR-100.
Method CIFAR-10 CIFAR-100
DN121 RN50 DN121 RN50
pgd-at 47.56 46.22 23.13 23.48
clat 49.91 49.45 25.74 25.81
clat (nd) 47.12 46.08 22.26 22.91
4.3.2 Ablating on critical layer selection
The choice of critical layer selection is another important feature impacting the performance of CLAT. We begin by examining the effect of dynamic layer selection. Table 6 and Table 6 highlight that dynamic selection is crucial to CLAT’s performance. Using the same layers throughout the process tends to cause overfitting and results in lower accuracies compared to the PGD-AT baseline.
To verify the significance of the selected critical layers, we compare CLAT with an alternative approach in which random layers are dynamically selected for fine-tuning instead of the critical layers. The results of this comparison are detailed in Table 7 and Table 12 (in Appendix C).
The data demonstrates that selecting critical layers significantly enhances the model’s adversarial robustness and clean accuracy. This observation is bolstered by our ablation study in Appendix Table 20 illustrating the performance effect of choosing the smallest versus largest critical indices for fine-tuning. Furthermore, Table 9 indicates near-identical critical layer selections within the same model, even across diverse datasets. This evidence supports our assertion that the variation in layer criticality arises from inherent properties within the model architecture, where certain layers are predisposed to learning non-robust features.
Table 7: Ablating CLAT Layer choices on CIFAR-10: The columns present PGD-10 and Auto Attack (AA) evaluation adversarial accuracies for models trained with CLAT, where layers are selected based on criticality or randomly.
Network Critical Layers Random Layers
Clean Acc. PGD-10 AA Clean Acc. PGD-10 AA
DN121 81.03 60.60 49.91 78.85 51.35 39.81
RN50 83.78 59.54 49.45 79.01 51.44 40.29
RN18 83.89 55.37 42.86 78.02 51.03 33.50
@@ 4.3.2 Ablating on critical layer selection
vgg19 39.3M 236K 0.6%
Lastly, we conduct an ablation study on the number of layers used in CLAT for fine-tuning. Fig. 3 and Fig. 7 illustrate the trade-offs between adversarial accuracy and the number of layers selected, as well as clean accuracy and the number of layers, respectively. Interestingly, both adversarial and clean accuracy are optimized with the same number of layers. Initially, fine-tuning more layers enhances model performance by increasing flexibility; however, this eventually diminishes CLAT’s effectiveness, likely because attention is diverted to less crucial layers at the expense of more important ones. This pattern underscores the critical role of specific layers in network robustness and emphasizes the need for deeper research into the dynamics of individual layers. Furthermore, we highlight that although the number of layers chosen impacts the learned robustness, CLAT still achieves robustness gains over the early-stopping baseline (no fine-tuning) with up to 10% of the layers selected, demonstrating its stability under small variations in the number of selected layers.
Figure 3: Comparative analysis on CLAT performance/PGD-10 adversarial accuracy with respect to number of critical layers used during CLAT
4.4 Overhead and stability analysis
The cost of optimizing the CLAT training objective in Eq. 6 is similar to that of the standard adversarial training given its min-max formulation. Here, we show that the computational overhead for determining critical indices is negligible. We verify in Table 10 that criticality indices can be stably computed with a single training batch as small as 10, with top-ranking layers consistent with those achieved with a larger batch. We further conducted over 1000 runs for each network to randomly select the data used for criticality estimation, where we find remarkable consistency in computed critical layers, differing in less than 5% of cases, typically involving only one layer change among the top five critical layers. The stability means we can use a mere 0.0002% of the training data for criticality estimation every 10 epochs, which introduces a neglectable 0.4% additional time to the standard adversarial training process as measured in Table 10.
@@ 5 Conclusions
In this work, we introduce CLAT, an innovative adversarial training approach that addresses robust overfitting issues by fine-tuning only the critical layers vulnerable to adversarial perturbations. This method not only emphasizes layer-specific interventions for enhanced network robustness but also sheds light on the commonality in non-robust features captured by these layers, offering a targeted and effective defense strategy. Our results reveal that CLAT effectively selects less than 4% critical trainable variables to achieve significant improvements in clean accuracy and adversarial robustness across diverse network architectures and baseline adversarial training methods. We limit the scope of this work to improving the empirical robustness of the model by utilizing the critical layers. In this sense, open questions remain on why these specific layers become critical, whether they can be identified more effectively, and whether the issues can be resolved with architectural or training scheme changes. We leave a more theoretical understanding of these questions as future work.
@@ Appendix A Pseudocode of CLAT
7:         Compute $\mathcal{W}_{\epsilon}(F_{i})$ for all layers with Eq. 2
8:         Compute $\mathcal{C}_{f_{i}}$ for all layers with Eq. 3
@@
13:         Perturbation $\delta\leftarrow$ Eq. 5 inner maximization
14:         $\mathcal{L}_{C}(f_{\mathcal{S}})\leftarrow$ Eq. 5
15:         Weight update $w[\mathcal{S}]$ with Eq. 6
@@ Appendix B Additional experiment results
We further compare the CLAT model robustness with the robustness of the SAT model against white-box attacks of various strengths. As illustrated in  Fig. 4, though both models are only trained against an attack of one strength ($\epsilon$ = 0.03), the improved robustness of CLAT is consistent across the full spectrum of attack strengths. This shows that CLAT is not overfitting to the specific attack strength used in training.
@@ C.2 Additional results on layer selection
Table 12: Ablating CLAT layer choices on CIFAR-100: The columns present PGD-10 and AutoAttack (AA) evaluation adversarial accuracies for models trained with CLAT, where layers are selected based on criticality or randomly.
Network Critical Layers Random Layers
Clean Acc. PGD-10 AA Clean Acc. PGD-10 AA
DN121 58.79 33.23 25.74 50.45 25.48 20.21
RN50 61.88 36.23 25.81 52.20 25.51 20.45
RN18 50.98 28.41 20.45 43.15 20.24 15.89
Appendix D Rebuttal Experiments
We thank the constructive feedback from all reviewers. We consolidate all additional experimental results suggested by the reviewers in this appendix section. We will integrate these results and our discussions in the rebuttal text into the finalized version.
D.1 Full Training Curves
Figure 5: White-box PGD-10 adversarial accuracy (y-axis) on CIFAR-10 for WRN34-10 models trained with CLAT fine-tuning starting at Epoch 70 (red), CLAT from scratch (orange), and PGD-AT (blue). The learning rate decays to 0 by Epoch 150.
D.2 Reduced Learning Rate Performance
Please see Fig. 6.
Figure 6: White-box accuracies (y-axis) for WRN34-10 on CIFAR-10 for models trained with the original learning rate multiplied by 0.1 at Epoch 70, using CLAT (green) and PGD-AT (red), compared to normal PGD-AT learning rate (purple).
D.3 Critical Index Variation Over Time
Please see Table 13.
Table 13: Critical layers identified at different epochs for various networks.
Network Ep. 70 Ep. 80 Ep. 90
DN121 [39, 14, 1, 3, 88] [38, 1, 5, 88, 15] [1, 5, 88, 2, 15]
RN50 [34, 41, 48] [48, 3, 36] [36, 2, 40]
RN18 [11] [11] [4]
D.4 Clean Accuracy and Critical Layer Selection
Figure 7: Comparative analysis on CLAT performance clean accuracies with respect to the number of critical layers used during CLAT.
D.5 Ablation Study: Dynamic and Fixed Critical Indices
Please see Table 14.
Table 14: Comparison of CIFAR-100 Clean and Adversarial Accuracies (PGD-10) on Different Networks with Fixed and Dynamic Layers for CLAT.
Model Fixed Layers Dynamic Layers
DN121 Clean 53.21 58.79
Adv 39.45 44.12
RN50 Clean 55.72 61.88
Adv 31.89 36.23
WRN34-10 Clean 56.45 62.38
Adv 28.56 32.05
D.6 Performance on Other Datasets
D.6.1 Imagenette
Results in Table 15.
Table 15: Comparative performance of CLAT across various networks on Imagenette. Robustness is evaluated with white-box PGD-10 and Auto Attack.
Model Method Clean (%) PGD-10 (%) AA (%)
DN121 PGD-AT 83.40 61.78 51.23
PGD-AT + CLAT 86.91 65.45 54.82
WRN70-16 PGD-AT 90.20 67.96 58.91
PGD-AT + CLAT 93.52 72.39 61.43
RN50 PGD-AT 84.02 62.10 50.45
PGD-AT + CLAT 87.11 64.89 54.31
WRN34-10 PGD-AT 90.45 65.45 56.31
PGD-AT + CLAT 93.21 69.82 60.04
VGG19 PGD-AT 85.45 56.71 46.53
PGD-AT + CLAT 89.72 59.45 51.22
RN18 PGD-AT 83.01 60.04 49.01
PGD-AT + CLAT 86.42 62.91 51.23
D.6.2 ImageNet
Results in Table 16.
Table 16: Clean and Adversarial Accuracies (PGD-10) performance comparison on ImageNet.
Model Method Clean Adv
DN121 PGD-AT 63.25 32.56
PGD-AT + CLAT 66.10 35.48
RN50 PGD-AT 65.88 33.18
PGD-AT + CLAT 67.12 36.91
WRN34-10 PGD-AT 64.31 31.08
PGD-AT + CLAT 66.12 33.59
D.7 Robustness against Other Attacks
We report that performance on varying strengths of PGD-10 in Appendix B. Table 17 highlights the robustness against stronger white-box attacks, including those not limited to $\ell_{\infty}$-bounded constraints.
Table 17: Adversarial accuracies across various attacks on CIFAR-10, highlighting the difference between models trained with the baseline and those utilizing CLAT. Positive blue values indicate the improvement in performance achieved with CLAT.
Attack Method DN121 (%) WRN34-10 (%)
FAB PGD-AT 44.80 40.12
PGD-AT + CLAT +3.70 +5.03
StAdv PGD-AT 48.50 45.15
PGD-AT + CLAT +0.91 +1.89
Pixle PGD-AT 10.40 9.50
PGD-AT + CLAT +2.21 +1.90
PGD-$\ell_{2}$ ($\epsilon$ = 0.03) PGD-AT 61.79 60.25
PGD-AT + CLAT +1.92 +1.45
$\ell_{\infty}$ ($\epsilon$ = 0.03, 50 steps) PGD-AT 57.01 54.01
PGD-AT + CLAT +1.92 +2.03
$\ell_{\infty}$ ($\epsilon$ = 0.03, 100 steps) PGD-AT 56.89 53.12
PGD-AT + CLAT +1.63 +1.89
D.8 Additional Performance Comparisons to Baselines
D.8.1 Stochastic Weight Averaging and Adversarial Weight Perturbation
Table 18 shows results when CLAT is augmented with SWA (Hwang et al., 2020) and AWP (Wu et al., 2020a) techniques. Omitted values are not reported in the original work.
Table 18: PGD-10 Adversarial Accuracies on CIFAR-10 and CIFAR-100 for PreAct RN-18 and WRN34-10 compared to baselines.
Network Method CIFAR-10 CIFAR-100
PreAct RN-18 AWP (Wu et al., 2020a) 55.39 30.71
AWP + CLAT 58.41 33.97
SWAAT (Hwang et al., 2020) 58.32 28.43
SWAAT + CLAT 60.76 30.74
WRN34-10 AWP 58.10 -
AWP + CLAT 60.89 -
SWAAT 61.45 31.97
SWAAT + CLAT 63.82 34.55
D.8.2 Data Augmentation Techniques
Please see Table 19.
Table 19: Adversarial accuracies on CIFAR-10 for PreAct RN-18 (using PGD-10) and WRN70-16 (using Auto Attack).
Network Method Adv. Accuracy (%)
PreAct RN-18 Data Aug (weak) (Li & Spratling, 2023) 50.34
Data Aug (weak) + CLAT 54.01
Data Aug (strong) 49.99
Data Aug (strong) + CLAT 52.37
WRN70-16 Wang et al. (2023) 70.69
Wang et al. (2023) + CLAT 72.34
D.9 Timing Comparisons
For a fair comparison, we use the same GPU configuration and number of GPUs across all methods, as described in the methods section. For DN121, one epoch of PGD-AT takes 67 seconds, RiFT takes 56 seconds per epoch, CLAT takes 69 seconds per epoch, and AutoLoRA also takes 69 seconds per epoch. As performed in the original paper for optimal performance, RiFT models were adversarially trained for 110 epochs, each taking 67 seconds, followed by fine-tuning for an additional 10 epochs at 56 seconds per epoch. Consequently, the total training time for RiFT is 132 minutes, compared to 112 minutes for CLAT (70 epochs of adversarial training and 30 epochs of fine-tuning).
D.10 Ablation: Effect of choosing largest critical indices/ most critical layers
Please see Table 20.
Table 20: Ablation study of CLAT layer choices on CIFAR-10: The columns present Clean, PGD-10, and Auto Attack (AA) evaluation accuracies for models trained with CLAT, where layers with the largest and lowest values are selected. The optimal number of layers per network was chosen for both approaches. All settings are consistent with the results in Table 1
. Network Largest cidx Smallest cidx PGD-AT Clean PGD-10 AA clean PGD-10 AA clean PGD-10 AA DN121 81.03 60.60 49.91 80.50 59.25 48.81 80.05 58.15 47.56 RN50 83.78 59.54 49.45 82.30 57.01 47.10 81.38 56.35 46.22 RN18 83.89 55.37 42.86 82.56 54.01 40.91 81.46 53.63 40.48
Appendix E Curvature-based weakness measurement
The main paper defines feature weakness based on the feature variation under worst-case perturbation. However, due to the non-linear and non-convex nature of the neural network model, the weakness measurement may not be precise in more complicated model architectures with a mixed layer type, such as the Vision Transformer model. To this end, this section provides a more accurate curvature-based formulation on feature weakness, and shows how the proposed weakness metric is an approximation. We leave the utilization of the curvature-based weakness measurement on more complicated models as future work.
Let’s start by considering the feature perturbation function $G_{i}(\cdot)$, which is defined at the output of layer $i$ on inputs close to a clean data point $x$:
$G_{i}(z)=||F_{i}(z)-F_{i}(x)||_{2}^{2}.$ (7)
The worst-case curvature of the function $G_{i}$ at the neighborhood of $z=x$ can be estimated following the formulation by Moosavi-Dezfooli et al. (2019) as
where $x^{\prime}$ is a worst-case perturbation (adversarial attack) maximizing $G_{i}(z)$ in the vicinity of $x$, and $\nabla G_{i}(x)=0$ by definition given it is a minimum. Following the observation in Moosavi-Dezfooli et al. (2019), a higher curvature indicates the feature to be more non-robust to adversarial examples. We can therefore use the curvature formulation $\nu_{i}(x)$ under a fixed perturbation budget $||x^{\prime}-x||_{p}\leq\epsilon$ to estimate the layer non-robustness, or weakness.
As we use the feature weakness to derive both the layer criticality metric and the finetuning objective, having a gradient term in the layer weakness leads to the costly computation of higher-order gradients in the optimization. To avoid the high cost of computing higher-order gradients when optimizing with the curvature, the numerator in the curvature formulation can be further derived as
In practice, it is also difficult to explicitly instantiate $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ for a neural network. To this end, we simplify the formulation in Eq. 9 by assuming $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ as a uniform vector. This leads to our definition of the $\epsilon$-weakness of layer $i$’s feature as:
where $N_{i}$ denotes the dimensionality of the output features at layer $i$, therefore normalizing the weakness measurement of layers with different output sizes. The weakness measurement is proportional to the curvature estimation in Eq. 9. A higher weakness value indicates that the feature vector is more vulnerable to input perturbations. The functionality of cascading layers from 1 to $i$ affects the vulnerability of the hidden features, as described by this formulation.